Suppose you borrowed $15,000 at a rate of 8.5% and must repay it in 5 equal installments at the end of each of the next 5 years. How much would you still owe at the end of the second year, after you have made the second payment? Here we use an annual compounding. Hint: Amortization loan table.
Answer #1
Annual payment = [P x R x (1+R)^N]/[(1+R)^N-1]
Where,
P= Loan Amount
R= Interest rate per period  
N= Number of periods
= [ $15000x0.085 x (1+0.085)^5]/[(1+0.085)^5 -1]
= [ $1275( 1.085 )^5] / [(1.085 )^5 -1
=$3806.49
Year Beginning Balance Interest Payment Ending Balance
a b c=b*8.5% d e=b+c-d
1 $                15,000.00 $ 1,275.00 $ 3,806.49 $        12,468.51
2 $                12,468.51 $ 1,059.82 $ 3,806.49 $           9,721.84
3 $                   9,721.84 $     826.36 $ 3,806.49 $           6,741.71
4 $                   6,741.71 $     573.05 $ 3,806.49 $           3,508.27
Correct Answer =$9721.84